OpenStudy (alfie):

What's wrong with my integration? I've used integration by parts: \[\int x\log ({x^2} + 1)dx \to \frac{{{x^2}}}{2}\log ({x^2} + 1) + \int \frac{{2{x^3}}}{{2{x^2} + 2}}\] Then the second integral I used polynomial division and got: \[\frac{{{x^2}}}{2} - log(2 + {x^2} + c)\] Sticking everything together to get: \[\frac{{{x^2}}}{2}\log ({x^2} + 1) + \frac{{{x^2}}}{2} - log(2 + {x^2}) + c\]

OpenStudy (alfie):

wolfram doesn't seem to agree with me: http://www.wolframalpha.com/input/?i=integrate%20xlog%28x^2%2B1%29dx&t=mfftb01

OpenStudy (alfie):

The + c wasn't supposed to be inside the parenthesis, equation typo >.<

OpenStudy (amistre64):

x+3/(x^2+1) is simplier i think

OpenStudy (amistre64):

x^3, not x+3 ... fat fingers

OpenStudy (amistre64):

and should be a - integration x^3/.... i believe

OpenStudy (amistre64):

\[\int x\log ({x^2} + 1)dx = \frac{{{x^2}}}{2}\log ({x^2} + 1) - \int \frac{x^2}{2}*\frac{2x}{x^2+1}dx \] \[\int x\log ({x^2} + 1)dx = \frac{{{x^2}}}{2}\log ({x^2} + 1) - \int \frac{x^3}{x^2+1}dx \]

OpenStudy (alfie):

Yup, I failed at the second integration (I factored out and left the +2 there, silly me), sec, let me do it and type it down :)

OpenStudy (slaaibak):

It would have been easier on the eye if you first substituted x^2 + 1 = u

OpenStudy (alfie):

Okay, I did it, it's all right I failed at factoring out (ridiculous >.<) Thank you amistre.

OpenStudy (amistre64):

is that log natural or base 10?

OpenStudy (amistre64):

youre welcome :)